OpenStudy (anonymous):

Every Monday, Rajiv receives fifty percent off enchilada dinners that regularly cost twenty-eight dollars. Which of the following shows the amount Rajiv pays for the dinners if he tips the waiter seventeen percent of the regular price for the dinners? Assume no sales tax is added to his bill. fifteen dollars and seventy cents eighteen dollars and seventy-six cents twenty-eight dollars thirty-one dollars

OpenStudy (anonymous):

rajiv pays 28/2 + 0.17x28 =14+4.76 =18.76 percentage=100*18.76/28 =67%

OpenStudy (anonymous):

b

OpenStudy (anonymous):

28(.5) = 14 + .17(28) = $18.67 B